It rains 3 days a week. When it rains there is Thunderstorm.The probability of Thunderstorm falling on the ground is 1/42000.A Golfer plays golf on Saturday and Sunday. What is the probability of Thunderstorm falling on that Golfer?

Grade:10

1 Answers

Swapnil Saxena
12 years ago

The question can be done with help of Combination.

Total possibiliteis of days on which rain occurs = 7C3

Total possiblities containing Sunday or Saturday as one of the days = 7C3-5C3

Probiabiltiy that the thunderstorm occurs on the Sundays and Saturdays = (7C3-5C3)/7C3

Probablitlity tha thuder storms occur on that day = (7C3-5C3)/7C3 * 1/42000= 25/35*1/42000 = 5/294000